OpenStudy (anonymous):

Can someone help me with linear algebra. How do I evaluate this (5-5i)^6

OpenStudy (anonymous):

factor out 5, then expand using binomial theorem

OpenStudy (anonymous):

notice $$5-5i=5(1-i)=5\sqrt2\left(\frac{\sqrt2}2-\frac{\sqrt2}2i\right)=5\sqrt2\ \left(\cos\left(-\pi/4\right)+i\sin\left(-\pi/4\right)\right)$$

OpenStudy (anonymous):

its linear algebra. I have to use de moivres formula

OpenStudy (anonymous):

recall de Moivre's theorem which tells us \((\cos t+i\sin t)^n=\cos(nt)+i\sin(nt)\)

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

so$$(5-5i)^6=(5\sqrt2\ (\cos(-\pi/4)+i\sin(-\pi/4)))^6=(5\sqrt2)^6(\cos(-\pi/4)+i\sin(-\pi/4))^6$$ and clearly \((5\sqrt2)^6=5^62^3=15625\times8=125000\) whereas $$(\cos(-\pi/4)+i\sin(-\pi/4))^6=\cos(-6\pi/4)+i\sin(-6\pi/4)=\cos(3\pi/2)-i\sin(3\pi/2)$$

OpenStudy (anonymous):

now we just recall \(\cos(3\pi/2)=0\) and \(\sin(3\pi/2)=-1\) so:$$(5-5i)^6=125000(0-i)=-125000i$$

OpenStudy (anonymous):

ok..my prof gives the answer as (5^6)(2^3)[cos(pi/2) + i sin (pi/2)]

OpenStudy (anonymous):

idk how to get the pi.2...i was getting ur answer

OpenStudy (kainui):

The key here is that this applies to vectors on the unit circle. So you have to normalize (5-5i) if you don't automatically recognize that this is just the vector, \[\frac{ \sqrt{2} }{ 2 }-\frac{ \sqrt{2} }{ 2 }i \] multiplied by a scalar multiple, you can use linear algebra, like the dot product. It might be necessary to do this if it's not a nice angle even multiple of pi.

OpenStudy (anonymous):

oops good catch @0213 I had two \(-1\) so it's actually \(+125000i\)

OpenStudy (kainui):

I can elaborate or go into more depth if that sort of lost you. But realize: cos(x)+isin(x) is just a vector on the unit circle, and you can think of cos(x)*1 as cos(x) being the length and 1 being the unit vector in the x-direction. Then you can think of sin(x) as being the length and i being the unit vector in the y-direction.

OpenStudy (anonymous):

See I looked up some videos on this and I get what the videos say and do....however, when I come to do the homework my prof gives me...Idk how to do it or I get the wrong answer. I didn't get the way he was teaching it in class, thats why i looked up the videos and i didn't get why he did some things.....could you explain some things to me and tell me when i should do them

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

Sure, so what exactly are some of the things that you're unsure of when to do them/ what to do? I'm not really sure where to begin, help me help you. =)

OpenStudy (anonymous):

the goal is to write it in the form \(z=r(\cos\theta+i\sin\theta)\) so that we can use de Moivre's theorem i.e. \(z^n=r^n(\cos n\theta+i\sin n\theta)\)

OpenStudy (anonymous):

Ok. We just did two lessons so it won't be so long. So the norm? Is that just like the magnitude of the complex number?

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

the norm is the same thing as the magnitude of course :-)

OpenStudy (anonymous):

My prof also said that the argument has to be between 0 and 2pi

OpenStudy (anonymous):

but in the videos, they had it over ....what is the difference

OpenStudy (anonymous):

your goal is to write it in trigonometric form: